how much pure gold (24k) must be mixed with 60% gold to get 20 grams of 75% gold
Answer by mananth(16949)

......... percent ---------------- quantity
24k 100---------------- x gms
Gold 60% 60---------------- 20-x gms
Total 75.00% ---------------- 20
100x+60(20-x)=75*20
100x+1200-60x=1500
100x-60x =1500-1200
40x=300
/40
x=7.5gms